OpenStudy (anonymous):

prove \[4\sin^236-1=4\cos^272\]

OpenStudy (anonymous):

guys i got to do this question

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

18*2=36 18*5=90 36*2=72

OpenStudy (anonymous):

\[4\sin^236-1=-2(1-2\sin^236)+3\]

OpenStudy (anonymous):

\[=-2(\cos 72)+3\]

OpenStudy (anonymous):

guys original question is actually this \[\frac{1}{2}\sqrt{4\sin^236-1}=\cos^272\]

OpenStudy (anonymous):

i mean the original question is \[\frac{1}{2}\sqrt{4\sin^236-1}=\cos 72\]

OpenStudy (anonymous):

so maybe we shud prove \[4\sin^236-4\cos^272=1\]

OpenStudy (anonymous):

ie \[4(\sin^236-\cos^272)\]

OpenStudy (anonymous):

i think i need a good identity

OpenStudy (anonymous):

The double angle identity for cosine, maybe?

OpenStudy (unklerhaukus):

\[\cos 2u = 1-2\sin^2 u\] \[(\cos 2u)^2 = (1-2\sin^2 u)^2\]

OpenStudy (experimentx):

try solving quadratic equation for sin(36) ... you might be able to reduce it into well known identity

OpenStudy (experimentx):

and if you get the value of sin(36) then that's fine ... http://www.wolframalpha.com/input/?i=sin%2836%29

OpenStudy (anonymous):

okay so \[\cos^2 72=(1-\sin^236)^2=1-2\sin^236+\sin^436\]not good

OpenStudy (anonymous):

i'll be back ...thank you...

OpenStudy (experimentx):

what's the problem, it's quadratic in sin(36)^2 i suppose you get it's value = (5 - sqrt(5))/8 the value of sin(36) is well known

OpenStudy (anonymous):

\[1-2(\frac{5-\sqrt{5}}{5})+(\frac{5-\sqrt{5}}{5})^2\\=1-2(\frac{5-\sqrt{5}}{5})+1+\frac{2\sqrt{}5}{5}+\frac{1}{5}\]

OpenStudy (anonymous):

=1/5